i am not understandin the concept of dot product nd cross product...


plzz explain me ....i am not understanding WHY is it used for????thnku

Grade:10

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The Dot Product

We define the dot product of two vectors

     v = ai + bj  and w = ci + d

to be

          v . w = ac + bd

 

The Cross Product Between Two Vectors


                          Definition  
Let u = ai + bj + ck  and v = di + ej + fk be vectors then we define the cross product v x w by the determinant of the matrix:

                                

 

W e can compute this determinant as

       

        =  (bf - ce) i + (cd - af) j + (ae - bd) k  

 

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi